# Minchul Kim, Anil Jain, Xiaoming Liu

### v2

@@ 5.1.1 State-of-the-Art Performance
For 1:1 verification tasks, determining if two images belong to the same person, Verification Accuracy is a primary metric, indicating the overall percentage of correct matches and non-matches. On more challenging datasets involving template-based verification or significant variability, performance is frequently reported as the True Accept Rate (TAR) at a specific False Accept Rate (FAR), often FAR=0.01% or even lower, indicating the likelihood of correctly accepting a genuine match while maintaining a very low rate of false acceptances. In NIST’s terminology, it is equivalent to 1-FNMR at a fixed FMR. For 1:N identification tasks, where an image must be matched against a gallery of known individuals, Rank-k accuracy is often used. This measures the percentage of times the correct identity is found within the top k ranked matches (Rank-1 being the most stringent). For open-set identification tasks (where the probe identity may not be in the gallery), the True Positive Identification Rate (TPIR) at a given False Positive Identification Rate (FPIR) (e.g., FPIR=0.01%) is a standard metric, measuring the rate of correctly identifying individuals present in the gallery while controlling the rate of incorrectly matching unknown individuals to someone in the gallery.
@@
CFP-FP [27]: Current methods achieve extremely high verification accuracy, often exceeding 99%. Top performance is typically seen with models utilizing ViT [102] backbones (e.g., ViT-L, ViT-B, ViT-S variants) or deeper ResNet [16] architectures (e.g., ResNet-101, ResNet-200). Effective loss functions like AdaFace [22] and ArcFace [20] are prevalent among the leading methods. Furthermore, training on very large datasets like Glint360K [79], WebFace [55] is crucial for reaching the highest scores, with methods like TransFace [147] reporting accuracies above 99.3%.
@@
IJB-C [24]: IJB-C dataset presents a more challenging scenario involving template-based matching (comparing sets of images/video frames). Performance is often measured by the TAR@FAR=0.01%. SoTA methods, such as PFC [79] (utilizing ViT-L or ResNet200), KP-RPE [39] (with ViT-B), AdaFace [22], and TopoFR [152], achieve TAR values around 98% at 0.01%. Again, larger backbones (ViT-L [102], ResNet200 [16]) and extensive training data (WebFace42M [55], Glint360K [79]) are characteristic of the top-performing approaches.
@@
TinyFace [73] (Low-Resolution Recognition): TinyFace specifically addresses the difficulty of recognizing faces from very low-resolution images. As expected, performance metrics like Rank-1 identification accuracy are considerably lower than on high-resolution datasets. Leading methods, predominantly using ViT-B backbones combined with techniques like KP-RPE that make the model robust to misalignments and loss functions such as AdaFace that allow quality adaptive training achieve Rank-1 accuracies around 75-76%. Training on large datasets like WebFace12M is also helpful for performance. Methods like ARoFace [151] also show competitive results, highlighting the ongoing efforts to improve recognition under significant resolution constraints.
@@
IJB-S [25]: Similar to TinyFace, IJB-S contains low-quality imageries and presents faces extracted from surveillance footage. We report Surveillance-to-Still (S2S) protocol. Top performance, measured by S2S Rank-1 accuracy, reaches approximately 73%. Another characteristic of this dataset is that the template size is large, making it a suitable choice to evaluate the methods for template feature fusion methods [71, 143, 142].
@@ 5.2 Technology Evaluations by NIST
The National Institute of Standards and Technology (NIST) has been conducting independent evaluations of FR technologies since 1999. Initially under the Face Recognition Vendor Test (FRVT), these activities have expanded to include the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) [157]. NIST evaluations are critical for assessing the technology readiness of algorithms for real-world deployment in government and operational settings.
@@
Between 2014 and 2018, NIST reported that face recognition software improved by a factor of 20 in search accuracy [158], highlighting the rapid pace of advancement in the field. To date, NIST has evaluated over 400 algorithms [159]. While academic benchmarks are crucial for driving research, the NIST FRVT provides an ongoing, operational evaluation of both academic and commercial algorithms under various scenarios, serving as a key indicator of the absolute SoTA deployed in real-world systems. Fig. 15 plots recent FRVT 1:1 verification performance results, including the performance trajectory of a strong academic baseline (ResNet-200 trained on the large-scale WebFace42M dataset, highlighted in red) for comparison against numerous vendor submissions. The performance results from the FRVT evaluations are publicly available on the organizer’s website [159].
@@ 6 Current Challenges in Face Recognition
Figure 16: Illustration of performance trends of various face recognition benchmarks over time, illustrating eventual performance saturation as datasets become extensively explored. This saturation is an indication of the progress in the field. Currently low-quality imagery benchmarks such as TinyFace and IJB-S are far from being solved.
Despite the impressive advancements in face recognition performance, several critical challenges persist, especially when moving beyond controlled benchmark scenarios to real-world deployments. As illustrated in Fig. 16, existing benchmarks tend to reach saturation over time as methods advance and datasets become well-explored. This saturation, however, does not necessarily indicate that face recognition is a solved problem. Instead, it underscores the limitations of current evaluation methodologies in capturing the complexity of real-world conditions. Some challenging examples are shown in Fig 14. In cases involving low-quality images (e.g., IJB-S images), facial details are often absent, forcing models to rely on more visible soft biometric cues, such as beards or hair color, to make more informed guesses.
@@ 6.1 Recognition at Scale: Large Galleries
A significant limitation of current academic benchmarks is their inability to replicate the scale and complexity of practical biometric systems. Real-world FR deployments often involve galleries with millions to billions of identities, far exceeding the thousands typically evaluated in research benchmarks. For instance, India’s Aadhaar national biometric system contains biometric data—including facial images—of over 1.4 billion individuals. At such immense scales, even minor degradations in recognition accuracy translate into substantial absolute numbers of false identifications or misses, impacting millions of users.
To highlight the disparity between benchmark evaluations and real-world scenarios, an illustrative experiment was conducted using the IJB-S dataset. The baseline IJB-S gallery (202 identities) was augmented with external imposters ranging from 1,000 to 10,000 identities. Tab. VIII presents the performance degradation observed with increasing gallery sizes. The results demonstrate a clear decline in recognition accuracy as the gallery grows, underscoring the challenge posed by real-world scenarios involving large galleries and numerous imposters.
This decline clearly highlights that conventional benchmarks do not fully represent operational scenarios, particularly when dealing with large-scale galleries containing diverse, unstructured, and noisy identities.
@@
TABLE VIII: Performance degradation on IJB-S (Survillance to Single protocol) as the gallery size increases with imposters sampled from an external dataset.
@@
6.2 Multi-modal Recognition: Beyond Facial Imagery
As face recognition technologies move towards more challenging environments characterized by low resolution, extreme poses, occlusions, varying illumination conditions, and large-scale databases, reliance solely on facial imagery becomes increasingly insufficient. Real-world scenarios such as surveillance or public safety applications require robust identification techniques capable of handling severely degraded visual information.
To address these challenges, there is growing emphasis on integrating multiple biometric modalities. Incorporating additional cues such as body shape, gait, or even behavioral patterns significantly enhances recognition robustness. Traditionally outputs from multiple biometric modalites are combined using score fusion [160]. Score-level fusion combines similarity scores from multiple biometric modalities after similarity comparison. Common approaches include normalization methods like Z-score and min-max, likelihood ratio-based fusion, and simple aggregations such as mean, max, or min fusion [161, 162, 163, 164]. These techniques collectively improve robustness and accuracy in challenging recognition scenarios.
On the other hand, multi-modal biometrics can be conducted with the fusion at the input or feature level. SapiensID [41] proposes to combine face and body recognition under one model, offering particular promise in cross modality comparison, as body images offer larger visual area that can reliably distinguish individuals at lower image resolutions. Ultimately, the future of robust FR lies in embracing a multi-modal approach, harnessing complementary biometric modalities to overcome the inherent limitations of any single modality.
@@ 6.3 Capacity of Generative Models
DCFace [116], trained on 52k real face images, generates 20k new synthetic identities. In contrast, Vec2Face [122], trained on a much larger dataset (360k images), achieves up to 200k well-separated identities. This scaling behavior demonstrates that generative identity capacity is closely related to the diversity and richness of the real training data.
Recent work by Boddeti et al. [165] propose a principled statistical framework for estimating the upper bound of this capacity, framing it as a hyperspherical packing problem in the feature space of a face recognition model. They define capacity as the maximum number of identities that can be placed in this space without exceeding a predefined similarity threshold (related to a false acceptance rate). Their empirical estimates show that StyleGAN3 has a practical upper bound, approximately 1.43 million identities at a 0.1% FAR, which decreases sharply with stricter thresholds. For class-conditional models like DCFace, the capacity was significantly lower, due to its greater intra-class variation.
This raises a compelling question for the future: could synthetic datasets eventually surpass the utility of real datasets for training FR models? While current synthetic data often lags behind real data due to domain gaps and capacity limitations, the potential advantages of synthetic generation could be unparalleled control over attributes, scalability, and the ability to systematically generate data for rare conditions or underrepresented demographics [166].
Realizing this potential likely requires moving beyond current 2D generative paradigms. Integrating 3D modeling and rendering techniques stands out as a particularly promising direction. By leveraging explicit 3D representations, future generative pipelines could offer physically grounded control over geometry, pose, illumination, and material properties (like skin texture and reflectance), potentially generating synthetic faces with greater realism, diversity, and, crucially, more distinct and well-separated identities than achievable through purely data-driven 2D synthesis alone. DigiFace [115] explores this direction and the key limitation is in the domain gap. Further research exploring these hybrid approaches, alongside developing better methods to measure and maximize the effective identity capacity, will be key to determining if and how synthetic data can ultimately overcome the limitations of, and perhaps even outperform, real-world data collection for advancing face recognition.
TABLE IX: Performance Comparison of Foundation Models (FMs) in Face Recognition under Different Training Regimes. Accuracies are averaged over LFW, CALFW, CPLFW, CFP-FP, and AgeDB. Rank is 16 for LoRA. CosFace [19] is used to train the models.
@@ 6.4 Role of Foundation Models in Face Recognition
Foundation models (FMs) are large-scale models pretrained on extensive image or text datasets for general-purpose tasks, rather than task-specific objectives such as face recognition. These models provide both pretrained weights and robust feature representations derived from broad visual or textual domains. Chettaoui et al. [167] offer a comprehensive overview of the role of foundation models in FR. Their findings indicate that, since FR models are traditionally trained on large-scale datasets, the advantages of using FMs are not clearly observed at the large scale training data.
However, fine-tuning FMs in low-data settings can significantly improve their performance [167]. Key comparative results are shown in Tab. IX. However, obtaining large scale training dataset is not difficult for face recognition, the benefit of FMs is still to be probed. Future work should focus on identifying which fine-tuning techniques, such as LoRA [113], and which foundation models, like CLIP [168] or DINOv2 [169], offer the best starting points for FR applications. Additionally, there is a need to understand why the advantages of foundation models diminish when training with large-scale FR datasets.
Recently, LAFS [68] introduces pretraining on unlabeled face data using foundation models, effectively learning critical face recognition representations and achieving strong few-shot performance. This highlights the value of specialized pretraining and motivates further exploration of domain-specific self-supervised learning (SSL) for developing specialized FR foundation models. It also raises questions about their interaction with general-purpose foundation models and potential reasons why the benefits of general models may diminish on large-scale FR datasets.
@@ 6.5 Interpretability
Figure 17: Comparison of visualization methods for face-related tasks. Grad-CAM [170] highlights broad, less interpretable regions, whereas ECLoss [171] enforces semantic correspondence with activations on meaningful facial parts. Peak Activation with FAD [74] shows that activations become more spread out across the face with the application of FAD loss. Attention maps from PartFViT [172] demonstrate that ViT models concentrate on discriminative facial parts, e.g., eyes and nose.
@@
Deep learning models used for face recognition are frequently viewed as black boxes. Their internal decision-making processes, involving millions of parameters, are inherently opaque, making it challenging to understand precisely why a particular decision (match vs. no-match, high vs. low quality score) was reached. This lack of transparency hinders trust, complicates debugging, and makes it difficult to assess its confidence, and impossible to be presented as evidence in the court. Several interpretability and explainability techniques are being applied or explored in the context of face recognition:
@@
Saliency/Attribution Maps: Methods such as Grad-CAM [170] or SHAP [173] generate heatmaps highlighting the input image regions (pixels) most influential for the model’s decision. For Transformer-based FR models, analyzing internal attention weights can offer insights into which parts of the input representation the model focuses on during processing [174]. FAD [74] proposes spatial activation diversity loss to learn more structured face activation. Some examples are shown in Fig. 17.
@@
Concept-based Explanation: Moving beyond pixel importance, these approaches aim to link model decisions to higher-level, human-understandable concepts [175]. This could involve identifying the influence of specific facial attributes (e.g., eye shape, nose structure) or using methods like ECLoss [171] to directly explain learned features without extra annotations. Some examples are shown in Fig. 17.
@@
Counterfactual Explanation: These techniques explain a decision by showing minimal changes to the input that would alter the model’s output [176, 177] (e.g., “How would this face need to change to no longer match?”).
@@
Frequency-Domain Explanation: Another approach specifically investigated for FR involves analyzing the influence of different frequency components (e.g., low vs. high frequencies representing coarse structure vs. fine details) in the input images on the matching decision [178, 179]. This provides a different perspective beyond spatial explanations.
@@
LLM based Explanation: Recent advances demonstrate that large language models (LLMs) can significantly enhance the interpretability of FR systems. Traditional interpretability tools like saliency maps or concept attributions often highlight important facial regions but fall short of conveying a full narrative understandable to users or forensic analysts. Recent approaches leverage LLMs to bridge this gap: models like XAI-CLIP [180] and vision-language alignment methods [181] can generate natural language explanations describing why two face images are matched or not, citing attributes such as “similar eyebrow curvature, matching nose bridge width, and aligned mouth corners.” Concept bottleneck models [182, 183] explicitly model this process by learning interpretable features like glasses, facial hair before the recognition decision to be explained in terms of high-level, semantically meaningful traits rather than low-level pixel activations.

### v3

@@ 5.1.1 State-of-the-Art Performance
In 1:1 verification, Verification Accuracy is the proportion of correct matches and nonmatches. On difficult datasets, performance is given by the True Accept Rate (TAR) at a fixed False Accept Rate (FAR), usually FAR=0.01%, representing $1-\text{FNMR}$ at a fixed FMR. For 1:N identification, Rank-$k$ accuracy shows how often the correct identity is within the top $k$ results (Rank-1 is strictest). In open set identification, the True Positive Identification Rate (TPIR) at a given False Positive Identification Rate (FPIR), such as FPIR=0.01%, measures correct identifications while limiting false matches of unknowns.
@@
CFP-FP [cfpfp]: Current methods achieve extremely high verification accuracy, often exceeding 99%. Top performance is typically seen with models utilizing ViT [dosovitskiy2020image] backbones (e.g., ViT-L, ViT-B, ViT-S variants) or deeper ResNet [he2016deep] architectures (e.g., ResNet-101, ResNet-200). Effective loss functions like AdaFace [kim2022adaface] and ArcFace [deng2019arcface] are prevalent among the leading methods. Furthermore, training on very large datasets like Glint360K [an2021partial], WebFace [webface] is crucial for reaching the highest scores, with methods like TransFace [dan2023transface] reporting accuracies above 99.3%.
@@
IJB-C [ijbc]: IJB-C dataset presents a more challenging scenario involving template-based matching (comparing sets of images/video frames). Performance is often measured by the TAR@FAR=0.01%. SoTA methods, such as PFC [an2021partial] (utilizing ViT-L or ResNet200), KP-RPE [kim2024keypoint] (with ViT-B), AdaFace [kim2022adaface], and TopoFR [dan2024topofr], achieve TAR values around 98% at 0.01%. Again, larger backbones (ViT-L [dosovitskiy2020image], ResNet200 [he2016deep]) and extensive training data (WebFace42M [webface], Glint360K [an2021partial]) are characteristic of the top-performing approaches.
@@
TinyFace [cheng2018low] (Low-Resolution Recognition): TinyFace specifically addresses the difficulty of recognizing faces from very low-resolution images. As expected, performance metrics like Rank-1 identification accuracy are considerably lower than on high-resolution datasets. Leading methods, predominantly using ViT-B backbones combined with techniques like KP-RPE that make the model robust to misalignments and loss functions such as AdaFace that allow quality adaptive training achieve Rank-1 accuracies around 75-76%. Training on large datasets like WebFace12M is also helpful for performance. Methods like ARoFace [saadabadi2024aroface] also show competitive results, highlighting the ongoing efforts to improve recognition under significant resolution constraints.
@@
IJB-S [ijbs]: Similar to TinyFace, IJB-S contains low-quality imageries and presents faces extracted from surveillance footage. We report Surveillance-to-Still (S2S) protocol. Top performance, measured by S2S Rank-1 accuracy, reaches approximately 73%. Another characteristic of this dataset is that the template size is large, making it a suitable choice to evaluate the methods for template feature fusion methods [kim2022cluster, jawade2024proxyfusion, jawade2023conan].
@@ 5.2 Technology Evaluations by NIST
NIST has been conducting independent evaluations of FR technologies since 1999. Initially under the Face Recognition Vendor Test (FRVT), these activities have expanded to include the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) [nist_frte_fate]. NIST evaluations are critical for assessing the readiness of algorithms for real-world deployment.
@@
Between 2014 and 2018, NIST reported that FR software improved by a factor of 20 in search accuracy [nist_2018_progress], highlighting the rapid pace of advancement in the field. To date, NIST has evaluated over 400 algorithms [grother2019face]. While academic benchmarks are crucial for driving research, the NIST FRVT provides an ongoing, operational evaluation of both academic and commercial algorithms under various scenarios, serving as a key indicator of the absolute SoTA deployed in real-world systems. Fig. 14 plots recent FRVT 1:1 verification performance results, including the performance trajectory of a strong academic baseline (ResNet-200 trained on the large-scale WebFace42M dataset, highlighted in red) for comparison against numerous vendor submissions. The performance results from the FRVT evaluations are publicly available on the organizer’s website [grother2019face].
@@ 6 Current Challenges in Face Recognition
Although FR performance has advanced greatly, challenges persist in real-world use. As Fig. 15 shows, benchmark saturation reflects limits in current evaluations rather than problem resolution. Fig. 16 illustrates that for low-quality images (e.g., IJB-S), missing facial details force models to depend on soft biometric cues like beards or hair color.
Figure 15: Illustration of performance trends of various FR benchmarks over time, illustrating eventual performance saturation as datasets become extensively explored. This saturation is an indication of the progress in the field.
@@ 6.1 Recognition at Scale: Large Galleries
A major limitation of current academic benchmarks is their failure to match the scale and complexity of real biometric systems. Practical FR deployments often manage galleries with millions to billions of identities, far beyond the thousands used in research. For example, India’s Aadhaar system stores biometric data, including facial images, for over 1.4 billion people. At such scales, even small drops in recognition accuracy lead to large numbers of false matches or misses, affecting millions of users.
An experiment on the IJB S dataset was conducted to illustrate the gap between benchmark evaluations and real world scenarios. The baseline IJB S gallery (202 identities) was expanded with external imposters ranging from 1,000 to 10,000 identities. Tab. VII shows a performance drop as gallery size increases, revealing reduced recognition accuracy with larger galleries. This demonstrates that benchmarks can fail to represent real world conditions involving extensive galleries with diverse, unstructured, and noisy identities.
6.2 Practical Applications
While benchmark performance provides a useful measure of algorithmic progress, the ultimate test of FR lies in its deployment across real-world applications. Each application domain introduces unique constraints, ethical considerations and requirements that continue to shape FR research.
Surveillance and Security. Surveillance applications revealed limits of low-quality, cross-resolution FR, prompting benchmarks like IJB-S [ijbs] and TinyFace [cheng2018low] and quality-aware models [meng2021magface, kim2022adaface]. More recent BRIAR project [BRIAR] extends evaluation up to 1,000 m standoff distances, integrating body and gait cues for robust recognition [jager2025expanding]. Under such setting, face-only systems achieve under 80% TAR@1% FAR and about 84% Rank-20 accuracy, while multimodal fusion exceeds 90% TAR, underscoring the need for cues beyond the face for long-range recognition [liu2024farsight].
Cross Age and Child Recognition. Age progression causes strong intra-class variation. NIST studies [grother2019face] show low errors for middle aged adults but higher rates for the youngest and oldest. YFA dataset [bahmani2022yfa] shows similarity degradation with even short age gaps, for ages under 36 months. Child recognition is reliable only over short term, Systems should use age aware thresholds, periodic re-enrollment, and multimodal fusion.
Other Specialized Domains. In mobile authentication and access control, FR achieves high reliability with emphasis on real time performance and privacy on edge devices [patel2016continuous]. In forensic and post mortem identification, robustness to extreme degradations, aging, and cross spectral imagery remains active, amid debates on accuracy, bias, and evidentiary standards [zeinstra2018forensic]. These applications illustrate the breadth of FR but merit separate detailed study.
Figure 16: Examples of faces across different datasets: IJB-S [ijbs], TinyFace [cheng2018low], IJB-C [ijbc], and CFP-FP [cfpfp]. The top row shows gallery images, the middle row shows corresponding probe images, and the bottom row reports cosine similarity scores. Higher similarity scores (in green) indicate successful matches, while lower scores (in red) indicate mismatches. Features are extracted by KP-RPE [kim2024keypoint] trained on WF4M [webface].
6.3 Multi-modal Recognition: Beyond Facial Imagery
As FR technologies move towards more challenging environments characterized by low resolution, extreme poses, occlusions, varying illumination conditions, and large-scale databases, reliance solely on facial imagery becomes increasingly insufficient. Real-world scenarios such as surveillance or public safety applications require robust identification techniques capable of handling severely degraded visual information.
To address these challenges, there is growing emphasis on integrating multiple biometric modalities. Incorporating additional cues such as body shape, gait, or even behavioral patterns significantly enhances recognition robustness. Traditionally outputs from multiple biometric modalites are combined using score fusion [singh2019comprehensive]. Score-level fusion combines similarity scores from multiple biometric modalities after similarity comparison. Common approaches include normalization methods like Z-score and min-max, likelihood ratio-based fusion, and simple aggregations such as mean, max, or min fusion [nandakumar2007likelihood, poh2011unified, poh2007improving, vatsa2007integrating]. These techniques collectively improve robustness and accuracy in challenging recognition scenarios.
On the other hand, multi-modal biometrics can be conducted with the fusion at the input or feature level. SapiensID [sapiensid] proposes to combine face and body recognition under one model, offering particular promise in cross modality comparison, as body images offer larger visual area that can distinguish individuals at lower image resolutions. The future of robust FR lies in embracing a multi-modal approach, harnessing complementary biometric modalities to overcome the limitations of any single modality.
@@ 6.3 Multi-modal Recognition: Beyond Facial Imagery
TABLE VII: Performance degradation on IJB-S (Survillance to Single protocol) as the gallery size increases with imposters sampled from an external dataset.
@@ 6.4 Capacity of Generative Models
DCFace [kim2023dcface], trained on 52k real face images, generates 20k new synthetic identities. In contrast, Vec2Face [wu2024vec2face], trained on a much larger dataset (360k images), achieves up to 200k well-separated identities. This scaling behavior demonstrates that generative identity capacity is closely related to the diversity and richness of the real training data.
Recent work by Boddeti et al. [boddeti2023biometric] propose a principled statistical framework for estimating the upper bound of this capacity, framing it as a hyperspherical packing problem in the feature space of a FR model. They define capacity as the maximum number of identities that can be placed in this space without exceeding a predefined similarity threshold (related to a false acceptance rate). Their empirical estimates show that StyleGAN3 has a practical upper bound, approximately 1.43 million identities at a 0.1% FAR, which decreases sharply with stricter thresholds. For class-conditional models like DCFace, the capacity was significantly lower, due to its greater intra-class variation.
TABLE VIII: Performance Comparison of Foundation Models (FMs) in FR under Different Training Regimes. Accuracies are averaged over LFW, CALFW, CPLFW, CFP-FP, and AgeDB. Rank is 16 for LoRA. CosFace [wang2018cosface] is used to train the models.
@@
This raises a compelling question for the future: could synthetic datasets eventually surpass the utility of real datasets for training FR models? While current synthetic data often lags behind real data due to domain gaps and capacity limitations, the potential advantages of synthetic generation could be unparalleled control over attributes, scalability, and the ability to systematically generate data for rare conditions or underrepresented demographics [um2024self].
Realizing this potential likely requires moving beyond current 2D generative paradigms. Integrating 3D modeling and rendering techniques stands out as a particularly promising direction. By leveraging explicit 3D representations, future generative pipelines could offer physically grounded control over geometry, pose, illumination, and material properties (like skin texture and reflectance), potentially generating synthetic faces with greater realism, diversity, and, crucially, more distinct and well-separated identities than achievable through purely data-driven 2D synthesis alone. DigiFace [bae2023digiface] explores this direction and the key limitation is in the domain gap. Further research exploring these hybrid approaches, alongside developing better methods to measure and maximize the effective identity capacity, will be key to determining if and how synthetic data can overcome the limitations of, and perhaps even outperform, real-world data collection for advancing FR.
@@ 6.5 Role of Foundation Models in Face Recognition
Foundation models (FMs) are large-scale models pretrained on extensive image or text datasets for general-purpose tasks, rather than task-specific objectives such as FR. These models provide both pretrained weights and robust feature representations derived from broad visual or textual domains. Chettaoui et al. [chettaoui2025froundation] offer a comprehensive overview of the role of foundation models in FR. Their findings indicate that, since FR models are traditionally trained on large-scale datasets, the advantages of using FMs are not clearly observed at the large scale training data.
However, fine-tuning FMs in low-data settings can significantly improve their performance [chettaoui2025froundation]. Key comparative results are shown in Tab. VIII. However, obtaining large scale training dataset is not difficult for FR, the benefit of FMs is still to be probed. Future work should focus on identifying which fine-tuning techniques, such as LoRA [hu2022lora], and which foundation models, like CLIP [radford2021learning] or DINOv2 [oquab2023dinov2], offer the best starting points for FR applications. Additionally, there is a need to understand why the advantages of foundation models diminish when training with large-scale FR datasets.
Recently, LAFS [sun2024lafs] introduces pretraining on unlabeled face data using foundation models, effectively learning critical FR representations and achieving strong few-shot performance. This highlights the value of specialized pretraining and motivates further exploration of domain-specific self-supervised learning (SSL) for developing specialized FR foundation models. It also raises questions about their interaction with general-purpose foundation models and potential reasons why the benefits of general models may diminish on large-scale FR datasets.
@@ 6.6 Interpretability
Figure 17: Comparison of visualization methods for face-related tasks. Grad-CAM [selvaraju2020grad] highlights broad, less interpretable regions, whereas ECLoss [Lin2024ECLoss] enforces semantic correspondence with activations on meaningful facial parts. Peak Activation with FAD [yin2019towards] shows that activations become more spread out across the face with the application of FAD loss. Attention maps from PartFViT [sun2022part] demonstrate that ViT models concentrate on discriminative facial parts, e.g., eyes and nose.
@@
Deep learning models used for FR are frequently viewed as black boxes. Their internal decision-making processes, involving millions of parameters, are inherently opaque, making it challenging to understand precisely why a particular decision (match vs. no-match, high vs. low quality score) was reached. This lack of transparency hinders trust, complicates debugging, and makes it difficult to assess its confidence, and impossible to be presented as evidence in the court. Several interpretability and explainability techniques are being applied or explored in the context of FR:
@@
Saliency/Attribution Maps: Methods such as Grad-CAM [selvaraju2020grad] or SHAP [lundberg2017unified] generate heatmaps highlighting the input image regions (pixels) most influential for the model’s decision. For Transformer-based FR models, analyzing internal attention weights can offer insights into which parts of the input representation the model focuses on during processing [rodis2024multimodal]. FAD [yin2019towards] proposes spatial activation diversity loss to learn more structured face activation. Some examples are shown in Fig. 17.
@@
Concept-based Explanation: Moving beyond pixel importance, these approaches aim to link model decisions to higher-level, human-understandable concepts [xai_concept]. This could involve identifying the influence of specific facial attributes (e.g., eye shape, nose structure) or using methods like ECLoss [Lin2024ECLoss] to directly explain learned features without extra annotations. Some examples are shown in Fig. 17.
@@
Counterfactual Explanation: These techniques explain a decision by showing minimal changes to the input that would alter the model’s output [stepin2021survey, xai_counterfactual] (e.g., “How would this face need to change to no longer match?”).
@@
Frequency-Domain Explanation: Another approach specifically investigated for FR involves analyzing the influence of different frequency components (e.g., low vs. high frequencies representing coarse structure vs. fine details) in the input images on the matching decision [Huber2024Frequency, Huber2025FrequencyBias]. This provides a different perspective beyond spatial explanations.
@@
LLM-based Explanation: Recent advances show that large language models (LLMs) are improving the interpretability of FR systems. Traditional tools such as saliency maps or concept attributions highlight key facial regions but fail to provide coherent, human-readable rationales. New approaches like XAI-CLIP [yao2024xaiclip] and interpretable vision–language alignment methods [liu2024interpretablevl] use LLMs to generate natural language explanations of why two faces match or differ, citing traits such as “similar eyebrow curvature, matching nose bridge width, and aligned mouth corners.” Complementary work on concept bottleneck models [koh2020concept, dombrowski2023faithful] enhances interpretability by learning high-level concepts (e.g., glasses, facial hair), allowing explanations grounded in semantic features rather than pixel activations. Such language-based reasoning enables FR systems to describe shared and divergent traits transparently, emphasizing the need for tighter grounding between textual justifications and visual evidence.